# CleAR: Robust Context-Guided Generative Lighting Estimation for Mobile Augmented Reality

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06submissionid: 9040ccs: Computing methodologies Mixed / augmented realityccs: Human-centered computing Ubiquitous and mobile computing systems and tools
@@ 1. Introduction
Related works on mobile AR lighting estimation systems seek to extract environment information from physical light probes (Prakash et al., 2019), user dynamics, and learning-based solutions (Zhao and Guo, 2021; Gardner et al., 2019; Wang et al., 2022; Yang et al., 2023). While physical light probes provide the most comprehensive environment observations, their use is limited by the need for physical setup. Consequently, AR applications often use learned models to estimate lighting from camera images. Over the past couple of decades, learning-based methods have evolved from discovering scene lighting cues from image details, such as highlights and shadows (Yu et al., 1999), to regressing omnidirectional environment lighting representations (Zhao and Guo, 2020, 2021; Gardner et al., 2019). However, autoregressive models cannot effectively tackle the environment information generation in lighting estimation. For example, Xihe (Zhao and Guo, 2021) provides real-time low-frequency lighting estimation for AR applications but lacks support for detailed environment reflections in object rendering. In contrast, recent advances in generative models enable the estimation of highly detailed environment maps (Wang et al., 2022; Yang et al., 2023), opening new possibilities for visually coherent rendering. Our work explores the integration of image-generative models into AR systems to enable high-quality environment lighting estimation, with novel AR context-aware design to improve generation speed, accuracy, and robustness. We summarize our main contribution as follows:
  • We design and implement CleAR, a novel AR context-guided generative lighting estimation system that overcomes the limited environmental sensing of AR devices. CleAR uses a two-step generative pipeline to produce accurate, visually detailed environment maps aligned with the user’s physical surroundings. We also create a large-scale lighting estimation dataset with $\sim$30K data items for training generative models.
  • We conduct a measurement study of recent lighting estimation datasets to analyze environmental lighting complexity and identify challenges to robustness. Guided by these insights, we design robustness testing experiments to evaluate several lighting estimation systems, including CleAR.
  • We develop a real-time refinement technique to enhance estimation quality. The refinement works by first matches the color appearance between the generated environment map and the real-time camera view to improve robustness, and then select the best environment map to match current lighting.
  • We implement CleAR as an end-to-end, edge-assisted framework that can be integrated into third-party AR applications. To demonstrate its usability, we integrate it into an example AR app for virtual object placement.
@@ 3. Motivation and Challenges
This section presents a measurement study of two key lighting attributes—light intensity and color temperature—to identify challenges for robust lighting estimation systems. Both attributes are essential for photorealistic and consistent rendering across diverse material types in mobile AR (Zhao et al., 2022a).
Our analysis uses two standard lighting estimation datasets: the Laval dataset (Gardner et al., 2017), an academic open research dataset, and the PolyHaven website (pol, 2024), a royalty-free HDR environment map data website. In total, we collected 2235 and 196 data items from the Laval dataset and PolyHaven, respectively. On these data, we calculate the light intensity as the total luminance of a given HDR environment map image, along with the environment map image’s color temperature values in Kelvin. Additional details of analysis setup, measurement calculations, and technical details are included in Appendix §A. Figure 2 shows the lighting condition distributions with annotated human-perceivable categories for light intensity and color temperature.
We make the following key observations. First, both datasets include a wide range of light intensity and color temperature values. In particular, the measured light intensity spans from as low as 0.6 to over 8 in luminance values, capturing both dim and brightly lit scenes. Similarly, the color temperature varies widely, ranging from below 1000K, typical of incandescent lighting, to over 9000K, which is seen in artificial lights. This suggests that real-world lighting conditions have complex and diverse properties. Second, there is a notable imbalance in the distribution of lighting conditions across the datasets. Specifically, both datasets are skewed toward neutral light intensity and color temperature values. The Laval dataset contains more low-light environment samples than bright-light ones. Meanwhile, a substantial portion of the PolyHaven dataset’s color temperature values fall within the cooler range.
@@
Our analysis reveals a critical challenge for achieving robust lighting estimation: complex and diverse real-world lighting conditions. In many real-world scenarios, environmental lighting conditions often vary across scenes and over time. For example, indoor rooms lit by sunlight often exhibit neutral to warm color temperatures, while those illuminated by LED sources tend to appear cooler. Therefore, to achieve robust estimation accuracy, sophisticated system designs are needed to understand and adapt to diverse environmental lighting conditions. In particular, lighting estimation systems must have rich built-in knowledge of diverse environmental lighting conditions, as well as adaptive application-time policies to react to complex AR device sensor outputs and user movements. In other words, lighting systems should be capable of learning generalizable lighting representations and not be overly sensitive to changes in environmental lighting. To address this challenge, in §4, we introduce an end-to-end system that takes multimodal AR data to guide the generation of a visually coherent HDR environment map with color refinement and color palette matching.
@@ 4.1. Overview
We design CleAR to support robust and high-quality lighting estimation for mobile AR. CleAR estimates omnidirectional environment lighting, a 360 HDR environment map image, from multimodal AR context data. In this work, we consider four types of AR data, including low-FoV LDR images, ambient light intensity and color temperature, and scene semantic map.
At a high level, estimating HDR omnidirectional lighting from low-FoV LDR environment observations is a process of significant information increase. To tackle this challenging problem, CleAR leverages the recent advancements in conditional image generative models (Zhang et al., 2023) to create missing environment lighting information, guided by AR context data. The key design of CleAR is a novel integration of image generative models to lighting estimation systems to ensure high-quality and robust estimation results.
As illustrated in Figure 3, the design of CleAR revolves around an AR context-guided generative lighting estimation pipeline (§4.2). While generative models excel at image generation in general, producing accurate HDR environment maps requires precise conditioning to ensure visual consistency with real-world scenes, especially under complex lighting conditions and limited environmental observations. We address this challenge by using AR context data to guide generation, offering both coarse- and fine-grained conditioning.
Another critical aspect of robust mobile AR lighting estimation is the timely update of estimation results to address changes in environmental lighting. However, the high inference latency of generative models limits the frequency of lighting updates. To address this challenge, CleAR employs an estimation refinement workflow (§4.3) that adapts the color profile of estimated environment maps to real-time AR camera images, ensuring temporally consistent lighting estimation. Specifically, CleAR first employs a multi-output generation strategy (§LABEL:subsubsec:multi_output_generation) to create reasonable yet diverse estimation results for unobserved environments. Then, CleAR uses several techniques (§4.3.1 and §4.3.2) to refine the previously estimated environment maps and select the best one as the estimation output. This workflow is well-suited for the edge-assisted AR setup, where computationally intensive generative models run on edge servers, while AR devices handle lightweight, real-time refinement.
@@ 4.3. Real-time Estimation Refinement
Completing 360 environment maps from low-FoV environment observations is an inherently ambiguous process. Although AR context data helps condition the generation process, each generation can take several seconds—leading to potential mismatches between the current AR camera image and the one used during generation. To address this, we propose a multi-output estimation process that enables real-time refinement. Our key insight is to generate multiple completion variants and select the one that best matches the real-time AR view. At a high level, the process consists of three steps as illustrated in Figure 5. First, CleAR uses different random seeds in the LDR completion network to output different completion results. Next, CleAR adjusts the color appearance of the LDR environment maps based on real-time AR camera image (§4.3.1). Finally, CleAR selects the optimal result that matches the color palette of the current AR view (§4.3.2).
@@
Our multi-output generation design not only enables CleAR to output accurate lighting estimation results, but also provides practical support for responsive lighting estimation in the real-time AR application workflow. Specifically, generative models, particularly diffusion-based models, typically suffer from long inference latency, making real-time estimation (¿30 FPS) impractical on mobile AR devices. Moreover, the high memory footprint of such models, often exceeding 2–4 GB during inference (Rombach et al., 2021), makes them unable to run on resource-constrained mobile platforms that must concurrently handle rendering and sensing tasks. In contrast, our refinement method enables the estimated environment maps to adapt to minor changes in environmental lighting, thereby reducing the need for frequent inference of the generative model. To facilitate real-time deployments, CleAR adopts an edge-device collaborative estimation architecture, similar to recent works (Ben Ali et al., 2022; Zhao and Guo, 2021). Specifically, LDR environment map completion and high-intensity map estimation can be offloaded to an edge server, while lightweight color refinement and result selection algorithms are executed on-device.
@@ 4.3.1. Color Appearance Adaptation
Addressing unpredictable changes in environmental lighting is crucial for supporting temporally consistent rendering of AR objects. Therefore, CleAR employs a lightweight on-device refinement technique to refine the multi-output LDR completion results to adapt their color appearances to real-time AR camera images. Our design selects the AR camera image as the adaptation target because the ultimate goal of mobile AR lighting estimation is to support the visually coherent insertion of AR objects into camera images. Additionally, our on-device real-time color appearance adaptation design is well-positioned to address the real-time estimation needs in mobile AR. Specifically, by applying color adaptation to the estimated environment maps, CleAR achieves consistently high estimation accuracy with few generative model inference requests. A recent work, DiffusionLight (Phongthawee et al., 2023), uses a similar strategy that has to invoke large quantities of inferences to achieve desirable quality results. However, as we will show in §6, DiffusionLight incurs a very high end-to-end inference time (359.9s), making its multi-output estimation strategy infeasible for real-time AR applications. In comparison, using our color adaptation technique, CleAR can achieve better estimation accuracy than DiffusionLight while using approximately 110X less time. Additional implementation details and testing examples of color refinement are included in Appendix §D.
@@
As the final step of the real-time estimation refinement process, CleAR selects the best environment map image from the previously refined LDR environment map images. In this process, CleAR searches for the best estimation candidate by identifying the highest cosine similarity between the color palettes of the observation and estimation environment map images. Specifically, the color palettes are selected using the K-means algorithm, which identifies the five most common colors for each environment map image. The total cosine similarity is calculated as the sum of per-color cosine similarity between color palettes. The selected LDR environment map will later be combined with its corresponding high-intensity map to form a complete HDR environment map for mobile AR applications.
@@ 6.1. Testbed-Based Evaluation
Table 1. System performance breakdown. (Left) We show the breakdown of system latency under the default generative estimation settings, which use five LDR completion outputs and do not use semantic maps in the input. In the left table, on-device operations are marked in gray, edge operations are marked in blue, and network operations are marked in green. (Right) We show the end-to-end system execution time (excluding network time) under various system configurations. Our system can generate an environment map as fast as 2,383 ms. Under the default system configuration, which is marked in bold font, the end-to-end execution time is 3,269 ms.
@@
Device data preparation 0.5 ($\pm$ 0.1)
AR Data Offloading 31 ($\pm$ 10.1)
LDR completion x5 1957 ($\pm$ 9.3)
LDR env. map retrieval 72 ($\pm$ 8.1)
Estimation refinement 3 ($\pm$ 0.2)
Refinement result synchronization 10 ($\pm$ 0.6)
High-intensity map estimation 1166 ($\pm$ 10.7)
High-intensity map retrieval 30 ($\pm$ 8.9)
@@ 6.2. Data-Driven Evaluation
Evaluation Metrics. Following prior works (Wang et al., 2022; Akimoto et al., 2022; Phongthawee et al., 2023; Gardner et al., 2019), we use the FID score (Seitzer, 2020) to measure the diversity of generated environment maps. We use three image-based metrics—scale-invariant Root Mean Square Error (si-RMSE) (Eigen et al., 2014), Angular Error (LeGendre et al., 2019), and RMSE—to evaluate lighting estimation accuracy on rendered virtual spheres. We follow  (Zhao et al., 2022a) in using RMSE to evaluate the accuracy of environment maps. We follow  (Phongthawee et al., 2023) to map the 0.1st and 99.9th value percentiles to 0 and 1 when calculating RMSE.
@@
In Figure 9, we show qualitative comparisons between CleAR and baseline methods in three scenes, representing the neutral, warm, and cool lighting conditions. Compared to other methods, CleAR allows more accurate overall color tones rendering and creates more diverse reflection details on the rendered virtual spheres. Figure 10 presents a quantitative comparison between CleAR and two SoTA methods (Phongthawee et al., 2023; Wang et al., 2022) that can output HDR environment maps in both rendering quality and estimation latency. Under the same environmental observations, CleAR achieves the lowest scale-invariant and normalized RMSE values across all three virtual sphere types. Noticeably, for the mirror sphere, the most challenging material of the three, CleAR achieves approximately $50\%$ reduction in the scale-invariant RMSE values. On the other hand, we observe that CleAR generates environment maps with slightly higher angular errors. Upon further inspection, we suspect that the diverse environment map image details potentially caused the higher angular error because the angular error metric is sensitive to pixel-wise differences. Furthermore, CleAR achieves the lowest lighting estimation latency by large margins. In particular, compared to DiffusionLight (Phongthawee et al., 2023), a diffusion model-based method, CleAR takes significantly less time (110X). This reduction in estimation latency is due to CleAR’s fewer model inferences and lower per-inference latency compared to DiffusionLight. This result suggests that CleAR can enable more responsive user experiences in mobile AR applications.

### v3

@@ Abstract.
ccs: Computing methodologies Mixed / augmented realityccs: Human-centered computing Ubiquitous and mobile computing systems and tools
@@ 1. Introduction
Related works on mobile AR lighting estimation systems seek to extract environment information from physical light probes (Prakash et al., 2019), user dynamics, and learning-based solutions (Zhao and Guo, 2021; Gardner et al., 2019; Wang et al., 2022; Yang et al., 2023). While physical light probes provide the most comprehensive environment observations, their use is limited by the need for physical setup. Consequently, AR applications often use learned models to estimate lighting from camera images. Over the past couple of decades, learning-based methods have evolved from discovering scene lighting cues from image details, such as highlights and shadows (Yu et al., 1999), to regressing omnidirectional environment lighting representations (Zhao and Guo, 2020, 2021; Gardner et al., 2019). However, autoregressive models cannot effectively tackle the environment information generation in lighting estimation. For example, Xihe (Zhao and Guo, 2021) provides real-time low-frequency lighting estimation for AR applications but lacks support for detailed environment reflections in object rendering. In contrast, recent advances in generative models enable the estimation of highly detailed environment maps (Wang et al., 2022; Yang et al., 2023), opening new possibilities for visually coherent rendering. Our work explores the integration of image-generative models into AR systems to enable high-quality environment lighting estimation, with novel AR context-aware design to improve generation speed, accuracy, and robustness.
We summarize our main contribution as follows:
  • We design and implement CleAR, a generative lighting estimation system that allows mobile AR applications to acquire environment lighting to support visually coherent AR experiences. The relevant research artifacts are at https://github.com/cake-lab/CleAR.
  • We introduce a two-step generative pipeline to produce more accurate and visually coherent environment maps aligned with the user’s physical surroundings. This design effectively overcomes the limited environmental sensing of AR devices and smartly leverages multi-modal AR context to achieve accurate estimation.
  • To further improve estimation accuracy and real-time robustness, we develop a set of refinement techniques that work in tadam with the generative estimation pipeline. These refinement components use a multi-output estimation strategy to address ambiguity and adapt the estimated environment color based on real-time AR device camera observations.
@@ 3. Motivation and Challenges
Our analysis uses two standard lighting estimation datasets: the Laval dataset (Gardner et al., 2017), an academic open research dataset, and the PolyHaven website (pol, 2024), a royalty-free HDR environment map data website. On these data, we calculate the light intensity as the total luminance of a given HDR environment map image, along with the environment map image’s color temperature values in Kelvin. Additional details of analysis setup, measurement calculations, and technical details are included in Appendix §A. Figure 2 shows the lighting condition distributions with annotated human-perceivable categories for light intensity and color temperature.
@@
Our analysis reveals two critical challenges. First, real-world lighting conditions are complex and diverse. Both datasets include a wide range of light intensity and color temperature values. In particular, the measured light intensity spans from as low as 0.6 to over 8 in luminance values, capturing both dim and brightly lit scenes. Similarly, the color temperature varies widely, ranging from below 1000K, typical of incandescent lighting, to over 9000K, which is seen in artificial lights. This suggests that real-world lighting conditions have complex and diverse properties. In many real-world scenarios, environmental lighting conditions often vary across scenes and over time. For example, indoor rooms lit by sunlight often exhibit neutral to warm color temperatures, while those illuminated by LED sources tend to appear cooler. Therefore, lighting estimation systems must have rich built-in knowledge of diverse environmental lighting conditions, as well as adaptive application-time policies to react to real-time environment lighting changes. To address this challenge, in §4, we introduce an end-to-end system that takes multimodal AR context to guide the generation of a visually coherent HDR environment map with color refinement and color palette matching.
@@ 4.1. Overview
At a high level, estimating HDR omnidirectional lighting from low-FoV LDR environment observations is a process of significant information increase. To tackle this challenging problem, CleAR leverages the recent advancements in conditional image generative models (Zhang et al., 2023) to create missing environment lighting information, guided by AR context data. The key design of CleAR is a novel integration of image generative models to mobile AR lighting estimation systems to ensure high-quality and robust estimation results.
As illustrated in Figure 3, the design of CleAR revolves around an AR context-guided generative lighting estimation pipeline (§4.2). CleAR estimates omnidirectional environment lighting, a 360 HDR environment map image, from multimodal AR context data. In this work, we consider four types of AR data, including low-FoV LDR images, ambient light intensity and color temperature, and scene semantic map. Our context-guided generation design addresses the key limitation of image generative models. While generative models excel at image generation in general, producing accurate HDR environment maps requires precise conditioning to ensure visual consistency with real-world scenes, especially under complex lighting conditions and limited environmental observations. We address this challenge by using AR context data to guide generation, offering both coarse- and fine-grained conditioning.
Another critical aspect of robust mobile AR lighting estimation is the timely update of estimation results to address changes in environmental lighting. However, the high inference latency of generative models limits the frequency of lighting updates. To address this challenge, CleAR employs an estimation refinement workflow (§4.3) that adapts the color profile of estimated environment maps to real-time AR camera images, ensuring temporally consistent lighting estimation. Specifically, CleAR first employs a multi-output generation strategy (§4.3.2) to create accurate yet diverse estimation results for unobserved environments. Then, CleAR uses several techniques (§4.3.1 and §4.3.2) to refine the previously estimated environment maps and select the best one as the estimation output.
@@ 4.3. Real-time Estimation Refinement
Completing 360 environment maps from low-FoV environment observations is an inherently ambiguous process. Although AR context data helps condition the generation process, each generation can take several seconds—leading to potential mismatches between the current AR camera image and the one used during generation. To address this, we introduce a multi-output estimation process co-designed with our two-step generation pipeline. Our key insight is to generate multiple completion variants and select the one that best matches the real-time AR view. At a high level, the process consists of three steps as illustrated in Figure 5.
@@
Our multi-output generation design not only enables CleAR to output accurate lighting estimation results, but also provides practical support for responsive lighting estimation in the real-time AR application workflow. The high memory footprint of generative models, often exceeding 2–4 GB during inference (Rombach et al., 2021), makes them unable to run on resource-constrained mobile platforms that must concurrently handle rendering and sensing tasks. In contrast, our on-device refinement method enables the estimated environment maps to adapt to minor changes in environmental lighting, thereby reducing the need for frequent inference of the generative model. To facilitate real-time deployments, CleAR adopts an edge-device collaborative estimation architecture, similar to recent works (Ben Ali et al., 2022; Zhao and Guo, 2021). Specifically, LDR environment map completion and high-intensity map estimation can be offloaded to an edge server, while lightweight color refinement and result selection algorithms are executed on-device.
@@ 4.3.1. Color Appearance Adaptation
Addressing unpredictable changes in environmental lighting is crucial for supporting temporally consistent rendering of AR objects. Therefore, CleAR employs a lightweight on-device refinement technique to refine the multi-output LDR completion results to adapt their color appearances to real-time AR camera images. While other additional metrics like semantic consistency or structural similarity are also important for environment map quality, we consider recovering these properties as orthogonal problems that can be addressed separately in real-time environment reconstruction systems (Theodorou et al., 2022).
Our design selects the AR camera image as the adaptation target because the ultimate goal of mobile AR lighting estimation is to support the visually coherent insertion of AR objects into camera images. Additionally, our on-device real-time color appearance adaptation design is well-positioned to address the real-time estimation needs in mobile AR. By applying color adaptation to the estimated environment maps, CleAR achieves consistently high estimation accuracy with few generative model inference requests.
A recent work, DiffusionLight (Phongthawee et al., 2023), uses a similar strategy that has to invoke large quantities of inferences to achieve desirable quality results. However, generative models, particularly diffusion-based models, typically suffer from long inference latency. Under the default setting, DiffusionLight requires 60 diffusion model inference calls to generate one estimation result. As we will show in §6, DiffusionLight incurs a very high end-to-end inference time (359.9s), making its multi-output estimation strategy infeasible for real-time AR applications. In comparison, using our color adaptation technique, CleAR can achieve better estimation accuracy than DiffusionLight while using approximately 110X less time. Additional implementation details and testing examples of color refinement are included in Appendix §D.
@@
As the final step of the real-time estimation refinement process, CleAR selects the best environment map image from the previously refined LDR environment map images. In this process, CleAR searches for the best estimation candidate by identifying the highest cosine similarity between the color palettes of the observation and estimation environment map images. The color palettes are selected using the K-means algorithm, which identifies the five most common colors for each environment map image. The total cosine similarity is calculated as the sum of per-color cosine similarity between color palettes. The selected LDR environment map will later be combined with its corresponding high-intensity map to form a complete HDR environment map for mobile AR applications.
@@ 6.1. Testbed-Based Evaluation
Table 1. System performance breakdown. (Left) We show the breakdown of system latency under the default generative estimation settings, which use five LDR completion outputs and do not use semantic maps in the input. In the left table, on-device operations are marked with D, edge operations are marked with E, and network operations are marked with N. (Right) We show the end-to-end system execution time (excluding network time) under various system configurations. Our system can generate an environment map as fast as 2,383 ms. Under the default system configuration, which is marked in bold font, the end-to-end execution time is 3,269 ms.
@@
D Device data preparation 0.5 ($\pm$ 0.1)
N AR Data Offloading 31 ($\pm$ 10.1)
E LDR completion x5 1957 ($\pm$ 9.3)
E LDR env. map retrieval 72 ($\pm$ 8.1)
D Estimation refinement 3 ($\pm$ 0.2)
E Refinement result synchronization 10 ($\pm$ 0.6)
N High-intensity map estimation 1166 ($\pm$ 10.7)
N High-intensity map retrieval 30 ($\pm$ 8.9)
@@ 6.2. Data-Driven Evaluation
Evaluation Metrics. Following prior works (Wang et al., 2022; Akimoto et al., 2022; Phongthawee et al., 2023; Gardner et al., 2019), we use the FID score (Seitzer, 2020) to measure the diversity of generated environment maps. For evaluating the accuracy of environment maps, we follow  (Zhao et al., 2022a) to use RMSE. We use three image-based metrics—scale-invariant Root Mean Square Error (si-RMSE) (Eigen et al., 2014), Angular Error (LeGendre et al., 2019), and RMSE—to evaluate lighting estimation accuracy on rendered virtual spheres. The si-RMSE and RMSE quantify pixel-wise differences, with si-RMSE being insensitive to global intensity scaling. Angular Error focuses on differences in pixel chromaticity, emphasizing the assessment of environment lighting color properties. We follow  (Phongthawee et al., 2023) to map the 0.1st and 99.9th value percentiles to 0 and 1 when calculating RMSE.
@@
In Figure 9, we show qualitative comparisons between CleAR and baseline methods in three scenes, representing the neutral, warm, and cool lighting conditions. Compared to other methods, CleAR allows more accurate overall color tones rendering and creates more diverse reflection details on the rendered virtual spheres. Figure 10 presents a quantitative comparison between CleAR and two SoTA methods (Phongthawee et al., 2023; Wang et al., 2022) that can output HDR environment maps in both rendering quality and estimation latency. Under the same environmental observations, CleAR achieves the lowest scale-invariant and normalized RMSE values across all three virtual sphere types. Noticeably, for the mirror sphere, the most challenging material of the three, CleAR achieves approximately $50\%$ reduction in the scale-invariant RMSE values. On the other hand, we observe that CleAR generates environment maps with slightly higher angular errors. Upon further inspection, we suspect that the diverse environment map image details potentially caused the higher angular error because the angular error metric is sensitive to pixel color differences. Furthermore, CleAR achieves the lowest lighting estimation latency by large margins. In particular, compared to DiffusionLight (Phongthawee et al., 2023), a diffusion model-based method, CleAR takes significantly less time (110X).
@@ 9. Conclusion
Acknowledgements.
We thank the anonymous reviewers for their constructive feedback. This work was supported in part by NSF Grants #2105564, #2236987, #2346133, #2350189, #2402383, and a VMware grant.